OpenStudy (anonymous):

Solve the equation. Round the answer to the nearest tenth. 2.8x = 5.1 A. 1.7 B. 1.8 C. 1.9 D. 18.2

OpenStudy (anonymous):

i got 1.9

OpenStudy (calculusxy):

When we divide 5.1/2.8, we get 1.821. Round 1.821 to the nearest tenths, you get 1.8.

OpenStudy (anonymous):

i got 1.9

OpenStudy (calculusxy):

Remember that is the number at the right is 4 or less, it stays the same. If the number on the right is 5 or more, it round up.

OpenStudy (anonymous):

oh ok i forgot about that

OpenStudy (mathstudent55):

\(2.8x = 5.1\) Since you want x by itself, and x is being multiplied by 2.8, you need to divide the left side by 2.8. The rule is that you need to do the same operation to both sides of an equation, so you need to divide both sides by 2.8. \(\dfrac{2.8x}{2.8} = \dfrac{5.1}{2.8} \) \(\dfrac{\cancel{2.8}x}{\cancel{2.8}} = \dfrac{5.1}{2.8} \) \(x = 1.8214\) Now we need to round off to the nearest tenth. The 8 is in the tenths place. Since the digit after the 8 is 2, which is lees than 5, the 8 remains as 8, so the answer is: \(x = 1.8\)

OpenStudy (anonymous):

omg thanks so much @mathstudent55

OpenStudy (mathstudent55):

You're welcome.